Examine the following steps in the solution to 5𝑥2 − 5 = 0

5(𝑥2 − 1) = 0

5(𝑥 − 1)(𝑥 + 1) = 0

𝑥 − 1 = 0

or

𝑥 + 1 = 0

𝑥 = 1

or

𝑥

= −1

What happened to 5? Explain.


1
Expert's answer
2021-11-16T11:42:53-0500

"5x^2\u22125=0"

This is the given equation

"5x^2−5=0\\5(x^2−1)=0\\(x^2−1)=\frac{0}{5}\\x^2−1=0\\(x+1)(x−1)=0\\x=1 \\or\\ x=−1"


5 got divided by 0